# Obstructions for quantitative measure equivalence between locally compact groups

### v1

@@ 1 Introduction
Two groups are orbit equivalent if and only if they are measure equivalent with common fundamental domains (see [Fur11] for a proof).
This paper deals with measure equivalence in the setting of locally compact groups, a precise definition in this context will be given in Section 2.2. Measure and orbit equivalences have already been studied for such groups, see e.g. [BFS13, CLM17, KKR21a, KKR21, Pau24, DLIT25], and it is worth mentioning that measure equivalence and orbit equivalence are the same notions among non-discrete, locally compact, second countable groups [KKR21, Theorem A]. Before stating our main results, we briefly review the discrete case
@@
  • Furman on lattices in higher rank semi-simple Lie groups [Fur99];
  • Kida on mapping class groups [Kid08, Kid10];
  • Guirardel and Horbez on $\mathrm{Out}(F_{N})$ [GH21];
  • Horbez and Huang on right-angled Artin groups [HH22, HH24];
  • Escalier and Horbez on graph products [EH24].
However, Ornstein and Weiss proved that any two free probability measure-preserving actions of infinite amenable groups are orbit equivalent [OW80]. In particular, measure equivalence is trivial for such groups. In the non-discrete setting, Koivisto, Kyed and Raum proved that the class of amenable, locally compact, second countable consists in three measure equivalence classes: compact groups, non-compact unimodular amenable groups and non-unimodular amenable groups [KKR21, Theorem B]. To counter this flexibility phenomenon, we impose restrictions on the functions arising from measure equivalence couplings, namely the cocycles.
@@
In [Sha04], Shalom was the first to study quantitative measure equivalence, more precisely the case of $\mathrm{L}^{\infty}$ measure equivalence, for amenable groups. His goal was to prove the quasi-isometry invariance of Betti numbers for nilpotent groups. We say that a measure equivalence coupling $(\Omega,X_{\Gamma},X_{\Lambda},\mu)$ is mutually cobounded if there exist finite subsets $F_{\Lambda}\subset\Lambda$ and $F_{\Gamma}\subset\Gamma$ such that $X_{\Gamma}\subset F_{\Lambda}\cdot X_{\Lambda}$ and $X_{\Lambda}\subset F_{\Gamma}\cdot X_{\Gamma}$.
Theorem 1.4 ([Sha04]).
@@
Historically, the notion of $\mathrm{L}^{p}$ measure equivalence, for $p\geq 1$, was introduced in [BFS13], and more generally $(\varphi,\psi)$-integrable measure equivalence was first defined in [DKLMT22] to study the weaker notion of $\mathrm{L}^{p}$ orbit equivalence for $p<1$. For weaker assumptions than $\mathrm{L}^{\infty}$, which geometric properties of amenable groups can be captured? In [Aus16], Austin first proved that integrably measure equivalent groups of polynomial growth have bi-Lipschitz equivalent asymptotic cones.
@@
For instance, $V_{\mathbb{Z}^{d}}(n)\simeq n^{d}$. More generally, a celebrated result of Gromov [Gro81] states that virtually nilpotent groups are exactly the groups with polynomial growth. Given non-trivial groups $F$ and $\Gamma$, we define the wreath product $F\wr\Gamma$ as
@@
In the appendix of [Aus16], Bowen proved an obstruction for integrable measure equivalence, using the notion of volume growth of finitely generated groups.
Theorem 1.5 ([Aus16, Theorem B.2]).
@@
Theorem 1.7 ([DLIT25, Theorem A.5]).
Let $G$ and $H$ be non-discrete unimodular locally compact compactly generated groups and let $p\in]0,1]$. If there exists an $(\mathrm{L}^{p},\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$, then
@@
Let $G$ and $H$ be non-discrete unimodular locally compact compactly generated groups and $\varphi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be an increasing and subadditive map. If there exists a $(\varphi,\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$, then
@@
We focus on unimodular groups because $\varphi$-integrability is undefined in the non-unimodular setting. Indeed, a definition of measure equivalence exists in the non-unimodular case (see e.g. [KKR21a, Definition 3.2]), but asks for the more general notion of non-singular actions. Another obstruction for an extension to non-unimodular groups will be the fact that isoperimetric profile, that we now deal with, does not contain any interesting information in the case of non-unimodular locally compact groups (we refer the reader to [Tes13, comments before Section 5.2]).
@@
first mentioned by Vershik [Ver73]. Finally a well-known inequality due to Coulhon and Saloff-Coste [CS93] relates $j_{1,\Gamma}$ and $V_{\Gamma}$:
@@
  • $j_{p,\Gamma}(n)\simeq\ln(n)$ for any polycyclic group $\Gamma$ with exponential growth [Pit95, Pit00];
  • $j_{1,F\wr\Gamma}(n)\simeq(\ln(n))^{\frac{1}{d}}$ with $F$ finite, and $\Gamma$ having polynomial growth of degree $d\geq 1$ [Ers03];
  • Brieussel and Zheng [BZ21, Theorem 1.1] give a large description of asymptotic behaviours for isoperimetric profiles. For any non-decreasing function $f\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ such that $t\mapsto\frac{t}{f(t)}$ is non-decreasing, there exists a finitely generated group $\Gamma$ with exponential volume growth having isoperimetric profile $j_{p,\Gamma}(n)\simeq\frac{\ln(n)}{f(\ln(n))}$.
@@
In Section 2.3, we will introduce isoperimetric profiles in the setting of compactly generated locally compact unimodular groups.
Simplifying the proofs in [DKLMT22] for a more natural adaptation to the locally compact setting, we get the two following extensions of Theorem 1.9 to locally compact groups.
@@
Let $G$ and $H$ be non-discrete unimodular locally compact compactly generated groups and $p\geq 1$. If there exists an $(\mathrm{L}^{p},\mathrm{L}^{0})$ measure equivalence coupling from $G$ to $H$, then
@@
Let $G$ and $H$ be non-discrete unimodular locally compact compactly generated groups and $\varphi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a non-decreasing map such that $t\mapsto t/\varphi(t)$ is non-decreasing. If there exists a $(\varphi,\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$, then
@@
Remark 1.10.
Delabie, Koivisto, Le Maître and Tessera [DKLMT22] also proved rigidity results for asymmetric notions of measure equivalence (measure subgroup, measure quotient, measure sub-quotient).
First, they find obstructions for quantitative versions, similar to Theorems 1.7 and 1.9. In the case of measure subgroups, these have been generalized to locally compact groups by the second-named author [Pau24]. In [DKLMT22], the authors need an assumption on the fundamental domains, "at most-$m$-to-one coupling", which becomes "coarsely $m$-to-one coupling" in [Pau24]. For instance, in the case of measure equivalence, these assumptions are implied by coboundedness. Thus we already knew from [Pau24] that Theorems B and C hold in the particular case of cobounded measure equivalence.
Secondly, in the same vein as Shalom’s Theorem 1.4, the following is proven in [DKLMT22]: the existence of an at most $m$-to-one $\mathrm{L}^{\infty}$ measure subgroup from $\Gamma$ to $\Lambda$ is equivalent to the existence of a regular embedding $\Gamma\to\Lambda$. As a consequence of the quantitative results of [DKLMT22], the isoperimetric profiles are monotonous under regular embeddings. Once again, these results have been generalized to locally compact groups in [Pau24].
Main ingredients for the proof of Theorems A, B and C.
In [DKLMT22], the authors crucially use the fact that any bijection between countable groups, or at least subsets of countable groups, is measure preserving. Indeed, the Haar measures are simply the counting measures in this context. These bijections are precisely the ones provided by the cocycles. This is the main issue when extending their main results for non-discrete locally compact groups since cocycles have no reason to provide measure-preserving maps between the groups.
In [DLIT25, Proposition A.1], the following is proven using cross-sections. Given a measure equivalence coupling between unimodular non-discrete locally groups, it is always possible to slightly modify the fundamental domains so that the cocycles satisfy the desired assumptions, namely providing Haar measure-preserving bijections between certain random measurable subsets of the groups.
Another hypothesis which will be useful for our proofs is ergodicity. To this end, we will use ideas from [KKR21, Proposition 2.17 (ii)].
Further details are provided in Section 3.
@@
Given positive integers $k,\ell\geq 1$ and polynomial growth groups $\Gamma$ and $\Lambda$ of degrees $k+\ell$ and $k$ respectively, the obstructions provided in [DKLMT22] imply the following: there is no $(\mathrm{L}^{p},\mathrm{L}^{0})$ integrable measure equivalence coupling from $\Gamma$ to $\Lambda$ for any $p>\frac{k}{k+\ell}$. The existence of an $(\mathrm{L}^{p},\mathrm{L}^{0})$ integrable measure equivalence coupling for any $p<\frac{k}{k+\ell}$ have recently been proved in [DLIT25, Theorem 1.6]. For $\Gamma=\mathbb{Z}^{k+\ell}$ and $\Lambda=\mathbb{Z}^{k}$, the result was already known [DKLMT22, Theorem 1.9].
Now the question is the existence of such a coupling for the threshold $p=\frac{k}{k+\ell}$. For more general groups, not necessarily of polynomial growth, we would like to know if the bounds of integrability given by Theorems 1.6 and 1.9 can be reached. The first-named author has recently answered this question [Cor25, Theorems A,B and D] in the discrete setting and the following statements provide extensions to locally compact groups.
@@
Let $G$ and $H$ be locally compact compactly generated groups. Assume that there exist a non-decreasing function $f_{G}$ and an increasing function $f_{H}$ satisfying $f_{G}(n)\simeq j_{1,G}(n)$, $f_{H}(n)\simeq j_{1,H}(n)$ and the following assumptions as $t\to+\infty$:
@@
Let $G$ and $H$ be locally compact compactly generated groups. Assume that there exist two increasing functions $f_{G}$ and $f_{H}$ satisfying $f_{G}(n)\simeq V_{G}(n)$, $f_{H}(n)\simeq V_{H}(n)$ and the following assumptions as $t\to+\infty$:

### v2

@@ 1 Introduction
Two groups are orbit equivalent if and only if they are measure equivalent with common fundamental domains (see [14] for a proof).
This paper deals with measure equivalence in the setting of locally compact groups, a precise definition in this context will be given in Section 2.2. Measure and orbit equivalences have already been studied for such groups, see e.g. [2, 4, 23, 22, 26, 10], and it is worth mentioning that measure equivalence and orbit equivalence are the same notions among non-discrete, locally compact, second countable groups [22, Theorem A]. Before stating our main results, we briefly review the discrete case
@@
  • Furman on lattices in higher rank semi-simple Lie groups [15];
  • Kida on mapping class groups [20, 21];
  • Guirardel and Horbez on $\mathrm{Out}(F_{N})$ [17];
  • Horbez and Huang on right-angled Artin groups [18, 19];
  • Escalier and Horbez on graph products [13].
However, Ornstein and Weiss proved that any two free probability measure-preserving actions of infinite amenable groups are orbit equivalent [25]. In particular, measure equivalence is trivial for such groups. In the non-discrete setting, Koivisto, Kyed and Raum proved that the class of amenable, locally compact, second countable consists in three measure equivalence classes: compact groups, non-compact unimodular amenable groups and non-unimodular amenable groups [22, Theorem B]. To counter this flexibility phenomenon, we impose restrictions on the functions arising from measure equivalence couplings, namely the cocycles.
@@
In [29], Shalom was the first to study quantitative measure equivalence, more precisely the case of $\mathrm{L}^{\infty}$ measure equivalence, for amenable groups. His goal was to prove the quasi-isometry invariance of Betti numbers for nilpotent groups. We say that a measure equivalence coupling $(\Omega,X_{\Gamma},X_{\Lambda},\mu)$ is mutually cobounded if there exist finite subsets $F_{\Lambda}\subset\Lambda$ and $F_{\Gamma}\subset\Gamma$ such that $X_{\Gamma}\subset F_{\Lambda}\cdot X_{\Lambda}$ and $X_{\Lambda}\subset F_{\Gamma}\cdot X_{\Gamma}$.
Theorem 1.4 ([29]).
@@
Historically, the notion of $\mathrm{L}^{p}$ measure equivalence, for $p\geq 1$, was introduced in [2], and more generally $(\varphi,\psi)$-integrable measure equivalence was first defined in [9] to study the weaker notion of $\mathrm{L}^{p}$ orbit equivalence for $p<1$. For weaker assumptions than $\mathrm{L}^{\infty}$, which geometric properties of amenable groups can be captured? In [1], Austin first proved that integrably measure equivalent groups of polynomial growth have bi-Lipschitz equivalent asymptotic cones.
@@
For instance, $V_{\mathbb{Z}^{d}}(n)\simeq n^{d}$. More generally, a celebrated result of Gromov [16] states that virtually nilpotent groups are exactly the groups with polynomial growth. Given non-trivial groups $F$ and $\Gamma$, we define the wreath product $F\wr\Gamma$ as
@@
In the appendix of [1], Bowen proved an obstruction for integrable measure equivalence, using the notion of volume growth of finitely generated groups.
Theorem 1.5 ([1, Theorem B.2]).
@@
Theorem 1.7 ([10, Theorem A.5]).
Let $G$ and $H$ be unimodular locally compact compactly generated groups and let $p\in]0,1]$. If there exists an $(\mathrm{L}^{p},\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$, then
@@
Let $G$ and $H$ be unimodular locally compact compactly generated groups and $\varphi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be an increasing and subadditive map. If there exists a $(\varphi,\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$, then
@@
We focus on unimodular groups because $\varphi$-integrability is undefined in the non-unimodular setting. Indeed, a definition of measure equivalence exists in the non-unimodular case (see e.g. [23, Definition 3.2]), but asks for the more general notion of non-singular actions. Another obstruction for an extension to non-unimodular groups will be the fact that isoperimetric profile, that we now deal with, does not contain any interesting information in the case of non-unimodular locally compact groups (we refer the reader to [31, comments before Section 5.2]).
@@
first mentioned by Vershik [32]. Finally a well-known inequality due to Coulhon and Saloff-Coste [8] relates $j_{1,\Gamma}$ and $V_{\Gamma}$:
@@
  • $j_{p,\Gamma}(n)\simeq\ln(n)$ for any polycyclic group $\Gamma$ with exponential growth [27, 28];
  • $j_{1,F\wr\Gamma}(n)\simeq(\ln(n))^{\frac{1}{d}}$ with $F$ finite, and $\Gamma$ having polynomial growth of degree $d\geq 1$ [12];
  • Brieussel and Zheng [3, Theorem 1.1] give a large description of asymptotic behaviours for isoperimetric profiles. For any non-decreasing function $f\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ such that $t\mapsto\frac{t}{f(t)}$ is non-decreasing, there exists a finitely generated group $\Gamma$ with exponential volume growth having isoperimetric profile $j_{p,\Gamma}(n)\simeq\frac{\ln(n)}{f(\ln(n))}$.
@@
Delabie, Koivisto, Le Maître and Tessera [9] also proved rigidity results for asymmetric notions of measure equivalence (measure subgroup, measure quotient, measure sub-quotient). In their works, they often use a crucial assumption on couplings, called being at most $m$-to-one. To keep it simply, we define this notion in the context of measure equivalence: a coupling from $\Gamma$ to $\Lambda$ is at most $m$-to-one if the fibers of $\alpha(.,x)\colon\Gamma\to\Lambda$ have cardinality at most $m$ for almost every $x\in X_{\Lambda}$.
In the same vein as Shalom’s Theorem 1.4, they proved the equivalence between the existence of an at most $m$-to-one $\mathrm{L}^{\infty}$ measure subgroup from $\Gamma$ to $\Lambda$, and the existence of a regular embedding $\Gamma\to\Lambda$. By a regular embedding, we mean a Lipschitz map $f\colon\Gamma\to\Lambda$ such that $\sup_{\lambda\in\Lambda}{|f^{-1}(\{\lambda\})|}$ is finite. As a stunning consequence of the quantitative results of [9], the isoperimetric profiles are monotonous under regular embeddings.
It is worth noticing that the “at most $m$-to-one” assumption is essential in their adaptation of Shalom’s theorem, since the bijections $\alpha(.,x)$, $x\in X_{\Lambda}$, provided by the cocycle of an at most $m$-to-one $\mathrm{L}^{\infty}$ measure subgroup are almost surely regular embedding, thus proving the existence of such maps from $\Gamma$ to $\Lambda$. However, for finer quantitative versions, the assumption seems avoidable. Although they could not get rid of it for rigidity results about quantitative measure sub-quotient, they manage to state Theorem 1.9 about quantitative measure equivalence without this assumption. We explain at the end of this introduction the basic ideas, and compare them with ours.
An extension of such quantitative results for asymmetric versions of measure equivalence was first brought by the second-named author [26], where the assumption “at most $m$-to-one coupling” became “coarsely $m$-to-one coupling”. However the author could not avoid it to generalize Theorem 1.9 in the locally compact setting. We are now able to state such a generalization without this assumption.
@@
Let $G$ and $H$ be unimodular locally compact compactly generated groups and $p\geq 1$. If there exists an $(\mathrm{L}^{p},\mathrm{L}^{0})$ measure equivalence coupling from $G$ to $H$, then
@@
Let $G$ and $H$ be unimodular locally compact compactly generated groups and $\varphi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a non-decreasing map such that $t\mapsto t/\varphi(t)$ is non-decreasing. If there exists a $(\varphi,\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$, then
@@
Given positive integers $k,\ell\geq 1$ and polynomial growth groups $\Gamma$ and $\Lambda$ of degrees $k+\ell$ and $k$ respectively, the obstructions provided in [9] imply the following: there is no $(\mathrm{L}^{p},\mathrm{L}^{0})$ integrable measure equivalence coupling from $\Gamma$ to $\Lambda$ for any $p>\frac{k}{k+\ell}$. The existence of an $(\mathrm{L}^{p},\mathrm{L}^{0})$ integrable measure equivalence coupling for any $p<\frac{k}{k+\ell}$ have recently been proved in [10, Theorem 1.6]. For $\Gamma=\mathbb{Z}^{k+\ell}$ and $\Lambda=\mathbb{Z}^{k}$, the result was already known [9, Theorem 1.9].
Now the question is the existence of such a coupling for the threshold $p=\frac{k}{k+\ell}$. For more general groups, not necessarily of polynomial growth, we would like to know if the bounds of integrability given by Theorems 1.6 and 1.9 can be reached. The first-named author has recently answered this question [7, Theorems A,B and D] in the discrete setting and the following statements provide extensions to locally compact groups.
@@
Let $G$ and $H$ be unimodular locally compact compactly generated groups. Assume that there exist a non-decreasing function $f_{G}$ and an increasing function $f_{H}$ satisfying $f_{G}(n)\simeq j_{1,G}(n)$, $f_{H}(n)\simeq j_{1,H}(n)$ and the following assumptions as $t\to+\infty$:
@@
Let $G$ and $H$ be unimodular locally compact compactly generated groups. Assume that there exist two increasing functions $f_{G}$ and $f_{H}$ satisfying $f_{G}(n)\simeq V_{G}(n)$, $f_{H}(n)\simeq V_{H}(n)$ and the following assumptions as $t\to+\infty$:
@@
Main ingredients for the extensions to the locally compact setting.
We finally record the obstacles encountered in our work, and briefly describe the main tools.
  • For the discrete setting, the proofs in [9] cannot be immediately adapted to a proof for the locally compact setting. To extend it, we had to find a slightly different approach that we now briefly describe.
    First of all, the proofs in the discrete setting are easier with the “at most $m$-to-one” assumption. Indeed, since Haar measures are simply counting measures in this context, it implies that the random maps $\alpha(.,x)\colon\Gamma\to\Lambda$ and $\beta(.,x)\colon\Lambda\to\Gamma$ are not far from being Haar-measure preserving up to a multiplicative constant (although they are not bijections). Then the strategy of the authors to get rid of this assumption is to average and thus prove in fact that this property of being at most $m$-to-one occurs with high probability, which is enough to get Theorem 1.9.
    Our approach now consists in using Haar-measure preserving maps arising from cocycles. The example of orbit equivalence, namely measure equivalence with $X_{\Gamma}=X_{\Lambda}$, is particularly enlightening. Indeed, a much easier proof of Theorem 1.9 can be found in this setting, using the fact that the cocycles provide Haar-measure preserving $\alpha(.,x)\colon\Gamma\to\Lambda$ and $\beta(.,x)\colon\Lambda\to\Gamma$, since these are bijections and Haar measures are counting measures. Coming back to the general case of measure equivalence, the analogous property is the following: if $x$ lies in the intersection $X_{\Gamma}\cap X_{\Lambda}$ between the fundamental domains (up to translation, we can assume that the intersection has positive measure), then the maps $\alpha(.,x)$ and $\beta(.,x)$ induce bijections (inverses of each other) between sets $R_{X_{\Gamma}\cap X_{\Lambda}}^{\Gamma}(x)$ and $R_{X_{\Gamma}\cap X_{\Lambda}}^{\Lambda}(x)$ defined by
    $R_{X_{\Gamma}\cap X_{\Lambda}}^{\Gamma}(x)\coloneq\{\gamma\in\Gamma\mid\gamma\cdot x\in X_{\Gamma}\cap X_{\Lambda}\}$
    and similarly for $R_{X_{\Gamma}\cap X_{\Lambda}}^{\Lambda}(x)$, where the symbol $\cdot$ refers to the induced actions $\Gamma\curvearrowright X_{\Lambda}$ and $\Lambda\curvearrowright X_{\Gamma}$. For the locally compact case, it is proved in [10, Proposition A.1], using cross-sections, that we can assume without loss of generality that the fundamental domains intersect themselves (with positive measure) and that the random bijections $\alpha(.,x)\colon R_{X_{G}\cap X_{H}}^{G}(x)\to R_{X_{G}\cap X_{H}}^{H}(x)$ and $\beta(.,x)\colon R_{X_{G}\cap X_{H}}^{H}(x)\to R_{X_{G}\cap X_{H}}^{G}(x)$ preserve the Haar measures for almost every $x\in X_{G}\cap X_{H}$.
    We can in fact extend this property to almost every $x\in X_{\Lambda}$: $\alpha(.,x)$ induces a Haar-measure preserving map from $R_{X_{\Gamma}\cap X_{\Lambda}}^{\Gamma}(x)$ to its image (and similarly for $\beta(.,x)$ on $R_{X_{\Gamma}\cap X_{\Lambda}}^{\Lambda}(x)$, for almost every $x\in X_{\Gamma}$). To this end, we require an ergodic coupling, and the desired property is a direct consequence of the fact that $R_{X_{\Gamma}\cap X_{\Lambda}}^{\Gamma}(x)$ (resp. $R_{X_{\Gamma}\cap X_{\Lambda}}^{\Lambda}(x)$) is nonempty for almost every $x\in X_{\Lambda}$ (resp. $x\in X_{\Gamma}$). Note that ergodicity is not a restrictive assumption in the discrete setting, and even in the locally compact setting, as we will verify using the ideas from [22, Proposition 2.17 (ii)]. Further details are provided in Section 3.
    We thus want a new proof in the discrete setting, that only relies on this property of the cocycles, and that can be adapted in the locally compact setting. To use the properties of cocycles, the challenging part is to prove that we do not lose information when reducing $\Gamma$ to $R_{X_{\Gamma}\cap X_{\Lambda}}^{\Gamma}(x)$, given $x\in X_{\Lambda}$, and $\Lambda$ to $R_{X_{\Gamma}\cap X_{\Lambda}}^{\Lambda}(x)$, given $x\in X_{\Gamma}$. To this end, we look at quantities on average, similarly to the strategies of Delabie, Koivisto, Le Maître and Tessera. As an illustration, given a map $\theta\colon\Gamma\to\mathbb{R}_{+}$, Fubini’s theorem implies that
    $\text{average}_{x\in X_{\Lambda}}{\int_{R_{X_{\Gamma}\cap X_{\Lambda}}^{\Gamma}(x)}{\theta(\gamma)\mathrm{d}\lambda_{\Gamma}(\gamma)}}=\text{constant}\times\int_{\Gamma}{\theta(\gamma)\mathrm{d}\lambda_{\Gamma}(\gamma)}$
    (we refer the reader to the last calculations in the proof of Proposition 5.4 for instance), where $\lambda_{\Gamma}$ refers to the Haar measure of $\Gamma$, which is simply the counting measure. We can thus replace $\int_{\Gamma}{\theta(\gamma)\mathrm{d}\lambda_{\Gamma}(\gamma)}$ by $\int_{R_{X_{\Gamma}\cap X_{\Lambda}}^{\Gamma}(x)}{\theta(\gamma)\mathrm{d}\lambda_{\Gamma}(\gamma)}$ for a huge amount of elements $x\in x_{\Gamma}$ (with a multiplicative error).
  • We finally move on to Theorems D and E. For finitely generated groups, we only need to check finitely many properties to prove that a cocycle is $\varphi$-integrable, namely the convergence of integrals indexed on a finite generating set. In the non-discrete setting, for compactly generated groups, not only do we have to check uncountably many properties: the convergence of integrals indexed on a compact generating set; but the supremum of these integrals over this set must also be finite (see Definition 2.4). Since the first-named author [7] crucially used the finiteness of a generating set in the discrete setting, the generalization to the case of locally compact compactly generated groups is not immediate and requires more tools, such as the techniques in [2, Appendix A.2].